Grade 12th passWave Motion

an earth satellite i in circular orbit at a distance of 38400 km from earth surface what is its peroiod of one revolution in days? take mass of earth M= 6*10^24 kg and radius 6400km... B. 27.5 days by formulas of y=Re+H and V=under root whole value of GM/r wanna full solution step by step plz....

Medical Student
8 Years agoGrade 12th pass

1 Answer

Aman
8 Years ago
Answer is 2.97By the formula of T^2=4(pi)^2 (h+R)^3 ÷ GmSo here h=38400kmM=6*10^24R=6400kmSo putting these valuesT^2=4*(3.14)^2*(6400+38400)^3÷6.67*10^-11*6*10^4So by solving thisAnswer is 2.97